2) A bag contains 9 white and 12 black balls. Find the probability of drawing 5 black balls.
3)A bag contains 2 red, 3 green, 5 blue, and 2 yellow balls. Find the probability that balls of all the colours are represented in a sample if four balls are selected at random.

Answer by stanbon(75887)

A bag contains 9 white and 12 black balls. Find the probability of drawing 5 black balls.
# of ways to draw the 5 black balls : 12C5 = 792
# of ways to draw 5 balls without restriction: 21C5 = 20349
Ans: 12C5/21C5 = 0.03892
----------------------------------------
3)A bag contains 2 red, 3 green, 5 blue, and 2 yellow balls. Find the probability that balls of all the colors are represented in a sample if four balls are selected at random.
# of ways to succeed: 2*3*5*2 = 60
# of 4-ball groups in 12: 12C4 = 495
Probability: 60/495 = 0.1212..
==========================================
cheers,
stan H.